OpenStudy (anonymous):

Wouldn't I use the permutation formula to solve this question: How many positive even integers with at most 3 digits can be created using 2,3,4,5,6, and 7??

OpenStudy (anonymous):

no

OpenStudy (anonymous):

you have three choices for the last digit

OpenStudy (anonymous):

then 5 choices for the tens place and 4 choices for the hundreds place

OpenStudy (anonymous):

but that would just be for three digit numbers, it says at least three digits, so there is more work to do

OpenStudy (anonymous):

oh no, it says "at most' three digits

OpenStudy (anonymous):

for three digit numbers it is \(3\times 5\times 4\) for two digit numbers it is \(3\times 5\) and for one digit numbers it is 3

OpenStudy (anonymous):

add up for your answer

OpenStudy (anonymous):

how come only 3 for the hundreds place?

OpenStudy (raden):

i think for this case, the repeat numbers is allowed

OpenStudy (raden):

so, for 3 digits number there are 6 * 6 * 3 ways for 2 digits number there are 6 * 3 ways for 1 digit number there are 3 ways

OpenStudy (anonymous):

k I understand. thx again!

OpenStudy (raden):

@satellite73 , please verify me :)